A diagnostic test has a probability 0.95 of giving a positive result when applied to a person suffering from a certain disease, and a probability0.10 of giving a (false) positive when applied to a non- sufferer. It is estimated that 0.5 % of the population are sufferers. Suppose that the test is now administered to a person about whom we have no relevant information relating to the disease (apart from the fact that he/she comes from this population). Calculate the following probabilities:

(a) that the test result will be positive.

(b) that, given a positive result, the person is a sufferer.

(c) that, given a negative result, the person is a non-sufferer.


1
Expert's answer
2022-04-23T11:06:22-0400

Let A - the person is suffering from a disease,

B - the test is positive,

Then "P(A)=0.005,~P(B|A)=0.95,~P(B|\lnot A)=0.1"


a) We can split probability of event B into two parts: probability of B when A occurred + probability of B when A did not occur:

"P(B)=P(B|A)P(A)+P(B|\lnot A)P(\lnot A)=\\
0.95\cdot0.005+0.1\cdot0.995=0.10425\approx10.4\%"


b) According to the Bayes' theorem:

"P(A|B)=\frac{P(B|A)P(A)}{P(B)}=\\
=\frac{0.95\cdot0.005}{0.10425}\approx0.0456=4.56\%"


c) According to the Bayes' theorem:

"P(\lnot A|\lnot B)=\frac{P(\lnot B|\lnot A)P(\lnot A)}{P(\lnot B)}"

"P(A)+P(\lnot A)=1," because events "A" and "\lnot A" are complementary

Thus:

"P(\lnot A)=1-P(A)=1-0.005=0.995\\
P(\lnot B)=1-P(B)=1-0.10425=0.89575"

Given that the event "\lnot A" happened, events "B" and "\lnot B" remain complementary:

"P(B|\lnot A)+P(\lnot B|\lnot A)=1\Rarr\\
\Rarr P(\lnot B|\lnot A)=1-P(B|\lnot A)=\\
=1-0.1=0.9"

"P(\lnot A|\lnot B)=\frac{P(\lnot B|\lnot A)P(\lnot A)}{P(\lnot B)}=\\
=\frac{0.9\cdot0.995}{0.89575}\approx0.9997=99.97\%"